(1 point) The Binomial Theorem. Let

p(x)=(2x−1)5=ax5+bx4+cx3+dx2+ex+f.p(x)=(2x−1)5=ax5+bx4+cx3+dx2+ex+f.

Then

(1 point) The Binomial Theorem. Let p(x) = (2x - 1)5 = ar” + bx4 + cr3 + de? + ex+f. Then , and
Answer #1

(12–18 = 4.627°C-1° +5€622%2C_v+$25 on? +S¢z¢=2000&vt secum=13*+5&&(2.2561€ S - 3 = 32x² -8ox 480x3 _uor² tlor - + 1 dl voor